OpenStudy (anonymous):

easy, but I can't remember how to solve: 3x/10=7/10

OpenStudy (anonymous):

actually dont... 3x/10=7/10, 3x/10(10/7)=7/10(10/7)

OpenStudy (anonymous):

now what did u get?

OpenStudy (anonymous):

multiply both sides by 10 .

OpenStudy (anonymous):

or u can multiply 3x X 10= 7 X10 cross products.

OpenStudy (anonymous):

\[\frac{ 3x }{ 10 }=\frac{ 7 }{10 }\] \[\frac{ 3x }{10 }\times 10=\frac{ 7 }{10 } \times 10\] 3x=7 now you can find x

OpenStudy (anonymous):

^^^^ that how you find x.

OpenStudy (anonymous):

not making sense to me. I don't remember algebra:(

OpenStudy (anonymous):

then 7/3... ok, the 10 cancels out anyway

OpenStudy (anonymous):

so x= 2.66

OpenStudy (anonymous):

I mean 2.33

OpenStudy (anonymous):

about 2.33

OpenStudy (anonymous):

2.34

OpenStudy (anonymous):

7/3 or 2.33